# A Set-Sequence Model for Time Series

### v1

@@ Abstract
In many financial prediction problems, the behavior of individual units (such as loans, bonds, or stocks) is influenced by observable unit-level factors and macroeconomic variables, as well as by latent cross-sectional effects. Traditional approaches attempt to capture these latent effects via handcrafted summary features. We propose a Set-Sequence model that eliminates the need for handcrafted features. The Set model first learns a shared cross-sectional summary at each period. The Sequence model then ingests the summary-augmented time series for each unit independently to predict its outcome. Both components are learned jointly over arbitrary sets sampled during training. Our approach harnesses the set nature of the cross-section and is computationally efficient, generating set summaries in linear time relative to the number of units. It is also flexible, allowing the use of existing sequence models and accommodating a variable number of units at inference. Empirical evaluations demonstrate that our Set-Sequence model significantly outperforms benchmarks on stock return prediction and mortgage behavior tasks. Code will be released.
@@ 1 Introduction
Important time series problems in finance and economics, such as estimating loan risk and forecasting stock returns, often involve predicting outcomes for a varying number of correlated units. For example, consider predicting the outcomes for a pool of 10,000 loans with 50 features per loan. A full joint model requires handling a total of 500,000 features to capture latent cross-sectional effects such as contagion [1, 2, 30]. This is practically intractable, so a common approach in the literature augments unit-level and macroeconomic feature series with manually crafted summary features intended to capture the latent cross-sectional effects (e.g. [22] employs lagged default rate for each zip code). Aside from requiring domain expertise, this approach does not scale well and does not reliably capture all relevant cross-sectional dynamics.
This paper proposes a Set-Sequence model for time-series, which learns the latent cross-sectional effects, eliminating the need to handcraft features. Figure 1 shows the architecture. Harnessing the exchangeability of units [^1] Exchangeability means the identity of individual units (e.g. stock ticker) is irrelevant – unit-level behavior is determined solely by observed features. Formally, for a problem of size $M$, the unit-level feature vectors $X^{i}$ are exchangeable: $P(X^{1},X^{2},\dots,X^{M})=P(X^{\pi(1)},X^{\pi(2)},\dots,X^{\pi(M)})$, for any permutation $\pi$. , the Set-Sequence model first computes an order-invariant set summary at each period, creating a time series of cross-sectional features. This series augments the time series for each unit separately, enabling the model to capture group dynamics without handcrafted features. Next, the prediction for each unit is made independently by a sequence model operating on this augmented time series containing both unit-level and cross-sectional features. This permits the use of any preferred sequence model (e.g. Transformer, RNN). Viewed through this lens, the Set-Sequence model makes simple architectural changes that have significant practical implications: it provides a way to extend a unit-specific model to include cross-sectional dependencies without needing to handcraft features, and the model allows inference over a variable number of units. Moreover, the compute scales linearly in the number of units. Section 3 provides details on our methodology.
@@
Figure 1: The Set-Sequence model: The Set model estimates a cross-sectional summary at each time, in linear complexity over the number of units. This augments the original features of each unit and the Sequence model consumes this augmented series to make predictions for each unit independently. Note that the Set model has a look-back of $L$ time periods, where $L\geq 1$ is a model parameter. The number of units $M$ and time periods $T$ may vary at inference.
[…]
@@
We next evaluate the Set-Sequence model on a stock portfolio prediction task using actual return data covering 36,600 US stocks. The model selects daily portfolio weights to maximize the portfolio’s annual Sharpe ratio (a widely used performance criterion). We compare against state-of-the-art sequence models and the leading task-specific CNN-Transformer of [11]. Over a 15-year out-of-sample period, the Set-Sequence model achieves between 22% and 42% higher Sharpe ratios compared to baseline models. These improvements are highly significant from an economic standpoint, and translate into substantial investment gains.
@@
Finally, we evaluate the Set-Sequence model on a large-scale US mortgage risk prediction task, comparing it to a suite of sequence models and the leading task-specific deep learning model of [22]. The task involves classifying the monthly mortgage state (current, 30/60/90 delinquency, foreclosure, prepayment) given its history. The data cover 5 million loan-month observations, and contain 52 dynamic loan-level and macro-economic features. Efficiently harnessing cross-sectional patterns, the Set-Sequence model achieves significantly higher AUCs than the benchmark models. The model delivers more accurate risk metrics, enabling more effective risk management.
Our work underscores the importance of exploiting exchangeability in high-dimensional time series and shows that the Set-Sequence model is an effective tool for tackling this. Our model: (1) Outperforms strong benchmarks on both synthetic and real tasks. It accommodates a variable number of units at inference, which is of significant practical interest. (2) Efficiently addresses large problems due to its linear scaling along the cross-sectional dimension. (3) Is interpretable with a low-dimensional set summary offering a window into the dependency structure across units.
@@ 2 Related Literature
Traditional approaches to modeling multiple time series have relied on, for example, Vector Autoregressive (VAR) models [26], which scale quadratically in the cross-sectional dimension. Linear factor models [4, 29, 27, 28, 3] offer an improvement, greatly reducing dimensionality compared to classical VAR approaches. While these linear models struggle with non-linear dependencies, deep learning approaches to factor modeling [32] have emerged to address non-linearities, though they still require retraining when new units are introduced.
Recent deep learning architectures, while powerful, fail to exploit cross-sectional exchangeability while maintaining computational efficiency. Transformer variants [31, 41, 18, 40, 19, 42, 35, 20], State Space Models such as S4 [10], H3 [7], and others [39, 34], convolution based models (LongConv [8], Hyena [21], and Multilayer Perceptrons [6, 37] all scale quadratically in the cross-sectional dimension. This often necessitates modeling each unit individually with shared parameters, relying on handcrafted features to capture joint dependencies as in [22].
Some approaches have attempted to capture cross-sectional exchangeability more directly. [23] combine a local recurrent neural network (RNN) model with a global Gaussian Copula model, but their simple covariance structure may not capture non-linear joint dependencies and has only been demonstrated for single-feature units. Other work explores capturing global dependencies through low-rank matrix factorization trained iteratively alongside forecasting models [24], but this also needs refitting when new units are introduced. There is also research that exploits set structures, introduced by [38, 16], though it focuses on set representations across time rather than in the cross-section [12]. We provide a more in-depth literature review in Appendix A.
@@ 3 Set-Sequence Model
Assume that $X^{1},\dots,X^{M}$ are the input units to the Set-Sequence layer, and $Y^{1},\dots,Y^{M}$ are the outputs of the layer. We use temporal chunks as many temporal patterns depend on several time steps to process. These inputs are processed in temporal chunks of size $L$, to allow the summary at the current period to use information from recent previous periods, producing feature embeddings via an embedding network, $\phi$. These embeddings are then fed through a set network $\rho$ to produce a permutation invariant, low dimensional, set summary $F_{t}$
$F_{t}=\rho\left(\frac{1}{M}\sum_{i=1}^{M}\phi\left(X_{(t-L,t)}^{i}\right)\right)\in\mathbb{R}^{r},$ (1)
@@
${Y}^{i}_{t+1}$ $=\operatorname{SeqLayer}(\tilde{X}^{i}_{(1,t)}),$ (3)
for all $i\in\{1,\dots M\}$, and $t\in\{1,\dots,T\}$.
@@
Other architectures are explored in Appendix B.
Implementation Details
Unless noted, we use LongConv [8] for the sequence layer, two-layer FFNs with dropout for $\phi$, $\rho$, and $\psi$, and chunk size L = 3. MHA-Seq differs only by using five attention heads. Full hyperparameters are in Appendix C.
4 Baselines and Task Selection
@@
Common multivariate-forecasting benchmarks—ETTh1, ETTm1, Exchange, and Traffic (see details in [41])—are ill-suited for our study: their unit counts are small enough that a full joint model remains tractable or they lack a clear exchangeable-unit structure. We therefore evaluate on three datasets—a synthetic contagion setting, an equity-portfolio regression task, and a mortgage-risk classification task—each defined over exchangeable units. They contain 1000, 500, and 2500 units with 4, 79, and 52 features per unit, giving cross-sectional dimensions of 4000, 39500, and 130000 that meaningfully stress models built for high-dimensional cross-sections.
@@
We benchmark against five widely adopted sequence architectures— Transformer [31], S4 [9], H3 [7], Hyena [21], and LongConv [8]. Together they span attention-, state-space-, and convolution-based methodologies, each of which our Set-Sequence layer can augment. We treat them as task-agnostic baselines on all three datasets. Because these models share a common backbone structure, we hold input/output dimensions, hidden size, learning-rate schedule, and other hyperparameters fixed, so the only variation lies in the sequence layer (and whether a Set module is present). This yields a fair comparison across models. Each task section also reports results for the strongest task-specific domain baseline. Hyperparameters for all models are in Appendix C.
@@ 5 Synthetic Task
This section examines the performance of the Set-Sequence model in a setting similar to default risk prediction for corporate and other loans, where contagion has been shown to play an important role [1, 2, 30]. The goal is to predict the transitions probabilities in the next time step for each of the units.
Task Description
We simulate 1000 exchangeable units, each with binary feature $x\in\{0,1\}$ over 100 time steps. Units transition between three states (state 3 is absorbing/default) with transition matrix proportional to:
@@
Default intensity evolves as $\lambda_{x,t+1}=\beta\lambda_{x,t}+\alpha N_{x,t}$, where $N_{x,t}$ is the fraction of type-$x$ units defaulting at time $t$. Parameters $\mu=0.001$, $\alpha=4$, and $\beta=0.5$ yield an average default rate of $\sim$1%. The model captures contagion via $\lambda_{t}=(\lambda_{0,t},\lambda_{1,t})$, where defaults increase future default risk. The core challenge in this synthetic setting is modeling joint dependencies among many units via the low-dimensional latent factor $\lambda_{t}$, itself a function of each unit’s state.
@@ 5.1 Comparison with Sequence Models
Table 1: Comparison of Set-Seq model with Sequence Models operating on one unit at a time (single) as well as the full cross-section (joint).
[…]
LongConv (Joint) 0.0369 0.681
S4 (Joint) 0.0375 0.676
H3 (Joint) 0.0401 0.675
Transformer (Joint) 0.0357 0.506
Hyena (Joint) 0.0358 0.702
LongConv (Single) 0.00182 0.757
S4 (Single) 0.00155 0.758
H3 (Single) 0.00170 0.751
Transformer (Single) 0.00156 0.758
Hyena (Single) 0.00169 0.760
Set-Seq (Ours) 0.00028 0.801
Model Ablations
We conduct an ablation study of the Set-Sequence model on the synthetic task, with results summarized in Table 2. We vary the cross-sectional modeling component (None, Set, or Multi-Head Attention) and the sequence length (1 or 50) to assess their impact on predictive accuracy and computational cost. The largest performance gain comes from adding a set model for the cross-sectional dimension, compared to modeling units independently. While replacing the set model with Multi-Head Attention (MHA) further improves accuracy, it incurs a $\times3$ increase in training time and a 3.5×3.5$\times$3.5 × increase in memory usage, due to the quadratic scaling of attention with the number of units.
@@
Seq. Model Set Model Seq. Len. KL(true|predicted) AUC Rel. Epoch Time Rel. Mem

### v2

@@ Abstract
Many prediction problems across science and engineering, especially in finance and economics, involve large cross-sections of individual time series, where each unit (e.g., a loan, stock, or customer) is driven by unit-level features and latent cross-sectional dynamics. While sequence models have advanced per-unit temporal prediction, capturing cross-sectional effects often still relies on hand-crafted summary features. We propose Set-Sequence, a model that learns cross-sectional structure directly, enhancing expressivity and eliminating manual feature engineering. At each time step, a permutation-invariant Set module summarizes the unit set; a Sequence module then models each unit’s dynamics conditioned on both its features and the learned summary. The architecture accommodates unaligned series, supports varying numbers of units at inference, integrates with standard sequence backbones (e.g., Transformers), and scales linearly in cross-sectional size. Across a synthetic contagion task and two large-scale real-world applications—equity portfolio optimization and loan risk prediction—Set-Sequence significantly outperforms strong baselines, delivering higher Sharpe ratios, improved AUCs, and interpretable cross-sectional summaries.
@@ 1 Introduction
Many key problems in finance—such as constructing stock portfolios or estimating risk for a pool of loans—involve predicting outcomes for large populations of correlated units. Similar problems appear in science and engineering where the individual units may be sensors, devices, or customers. A key characteristic of such problems is they involve a varying number of correlated units, with each unit represented by a time series of dynamic features. In short, they involve a set of sequences where the sequences share the time axis. This poses a two-dimensional challenge: capturing dependencies across the cross-sectional dimension of $M$ units, while modeling the temporal dynamics over $T$ time steps. Directly modeling the joint distribution of the $M$ units is intractable due to its size. For example, a portfolio of $M=10,000$ loans where each loan has $50$ features per time step, results in a (joint) time series with $50,000$ features per time step (Anenberg & Kung (2014); Azizpour et al. (2018)). To avoid this, existing approaches make predictions for each unit separately, while augmenting the unit’s own features with hand-crafted summaries for the cross-section; the model learns the temporal dynamics but the cross-sectional features are hand-designed. For example, Sadhwani et al. (2020) employs aggregate foreclosure rate as a summary feature across the loans. This approach requires domain expertise and is unlikely to capture all latent effects.
To address this two-dimensional challenge, we propose the Set-Sequence model, which decouples the modeling of cross-sectional and temporal dependencies by learning latent cross-sectional effects directly from data. The architecture consists of two components (see Figure 1). First, a Set module processes the permutation-invariant cross-section at each time step, leveraging unit exchangeability [^1] Exchangeability implies that unit identities (e.g., stock tickers) are irrelevant—behavior is determined solely by observed features. Formally, for a problem of size $M$, the unit-level feature vectors $X^{i}$ are exchangeable: $P(X^{1},\dots,X^{M})=P(X^{\pi(1)},\dots,X^{\pi(M)})$ for any permutation $\pi$. to compute an order-invariant summary of the population state. Second, this summary is concatenated to each unit’s features and passed to a Sequence module (e.g., Transformer or RNN) that models temporal dynamics. The Set-Sequence model fulfills three key desiderata: inference over a variable number of units, handling unaligned units with different start and end times, and integration with standard sequence models. This simple architecture yields significant practical benefits: it removes the need for handcrafted features and seamlessly incorporates cross-sectional dependencies into unit-level predictions.
@@
Figure 1: The Set-Sequence model. The Set model estimates a cross-sectional summary at each time, in linear complexity over the number of units. This augments the original features of each unit and the Sequence model consumes this augmented series to make predictions for each unit independently. Note that the Set model has a look-back of $L$ time periods, where $L\geq 1$ is a model parameter. The number of units $M$ and time periods $T$ may vary at inference.
Theoretical analysis highlights key properties of the Set-Sequence model with direct practical implications. First, a forward-pass complexity result shows that mean-based cross-sectional aggregation scales linearly in the number of units $M$, in contrast to the quadratic scaling of attention-based methods—enabling efficient training on large cross-sections. Second, under exchangeability, a polynomial sufficiency result guarantees that pooled monomials up to degree $k$ can approximate any continuous permutation-invariant cross-sectional statistic. Together, these results motivate the use of simple set summaries as a scalable and expressive default, capable of capturing rich latent structure without incurring prohibitive computational cost.
We first benchmark the performance and efficiency of the Set-Sequence model on a synthetic contagion task that mimics default prediction for a large pool of exchangeable loans. Defaults are driven by a shared latent factor $\lambda_{t}$, capturing contagion effects observed in corporate and consumer credit markets. The Set-Sequence model substantially outperforms state-of-the-art sequence models, achieving near-optimal performance across a wide range of pool sizes, with accuracy improving as coverage increases. It also exhibits strong interpretability: the learned set summaries closely track the true latent factor ($\rho=0.95$). Notably, the simple linear aggregation used in the Set module outperforms more complex multi-head attention layers, both in accuracy and efficiency.
@@
We next evaluate the model on a real-world stock portfolio construction task using return data for 36,600 U.S. equities. The goal is to select daily portfolio weights that maximize the annualized Sharpe ratio (a widely used performance criterion). We compare Set-Sequence against leading sequence models and the task-specific CNN-Transformer of Guijarro-Ordonez et al. (2022). Over a 15-year out-of-sample period, Set-Sequence achieves 22–42% higher Sharpe ratios than the strongest baselines. These improvements translate into economically meaningful investment gains.
@@
Finally, we apply Set-Sequence to a large-scale U.S. mortgage risk prediction task, using 5 million loan-month observations and 52 dynamic loan-level and macroeconomic features. The task involves classifying the next-month mortgage state (e.g., current, 30/60/90 days delinquent, foreclosure, prepayment). By effectively capturing latent cross-unit dependencies, Set-Sequence achieves significantly higher AUCs than both general-purpose sequence models and the domain-specific deep learning benchmark of Sadhwani et al. (2020). These gains translate into more accurate risk metrics and improved decision support for large-scale mortgage portfolio management.
This work highlights the value of exploiting exchangeability in high-dimensional time series and demonstrates that the Set-Sequence model offers a practical and effective solution. It: (1) Outperforms strong baselines on both synthetic and real-world tasks; (2) Scales efficiently to large cross-sections, with linear complexity in the number of units and support for a variable number of units at inference; and (3) Provides interpretability, with low-dimensional set summaries that capture and reveal latent cross-sectional structure. Together, these properties make Set-Sequence a compelling architecture for modern time-series prediction problems involving large populations of interacting units.
@@ 2 Related Literature
Traditional methods for modeling multiple time series include Vector Autoregressive (VAR) models Sims (1980), which scale quadratically in the cross-sectional dimension. Linear factor models Chamberlain & Rothschild (1983); Stock & Watson (2002a; 2011) reduce dimensionality but struggle with non-linear dependencies. Deep learning-based factor models Wang et al. (2019) address this limitation, yet require retraining when new units are introduced.
Recent deep learning architectures for time series, though powerful, are typically not designed to exploit cross-sectional exchangeability as they rely on the specific identities of the units and do not scale to arbitrary units at inference. Models like iTransformer Liu et al. (2024) and TimeMixer Wang et al. (2024a) focus on fixed variates or multiscale temporal patterns, respectively. Some approaches model cross-sectional exchangeability more directly, using Gaussian Copulas Salinas et al. (2019) or low-rank matrix factorization Sen et al. (2019), but these impose restrictive assumptions or require refitting when units change. Permutation-invariant architectures such as Deep Sets Zaheer et al. (2017) and Set Transformers Lee et al. (2019) model unordered collections effectively, but are designed for static sets and lack temporal structure. They can handle a set of sequences, but our setting contains additional structure: the sequences live on a shared time axis. Our work introduces a novel dynamic architecture that applies set-based modeling at each time step, enabling scalable and expressive handling of high-dimensional, time-indexed cross-sections.
Graph Neural Networks Scarselli et al. (2009) offer a flexible framework for modeling cross-sectional dependencies, especially when an explicit relational graph is available. However, in large-scale applications such as mortgage risk prediction—with hundreds of thousands of loans—inferring a reliable dependency graph is challenging and computationally costly, even if the graph is learned as part of the objective Wu et al. (2020b). In contrast, the Set-Sequence model sidesteps this requirement by learning latent cross-sectional structure directly from observed features.
General-purpose models—such as Transformer variants, State Space Models (S4, H3), and MLP-based architectures Vaswani et al. (2017); Gu et al. (2022); Fu et al. (2022); Chen et al. (2023)—scale quadratically in the number of units, often requiring per-unit modeling with shared weights and handcrafted features to capture joint effects Sadhwani et al. (2020). The Set-Sequence model instead learns cross-sectional dependencies directly, with linear scaling and no feature engineering.
We provide a more in-depth literature review in Appendix B.
@@ 3 Set-Sequence Model
Let $X^{1},\dots,X^{M}$ are the input units to the Set-Sequence layer, and $Y^{1},\dots,Y^{M}$ are the outputs of the layer. We use temporal chunks as many temporal patterns depend on several time steps to process. These inputs are processed in temporal chunks of size $L$, to allow the summary at the current period to use information from recent previous periods, producing feature embeddings via an embedding network, $\phi$. These embeddings are then fed through a set network $\rho$ to produce a permutation invariant, low dimensional, set summary $F_{t}$
$F_{t}=\rho\Big(\frac{1}{M}\sum_{i=1}^{M}\phi\big(X_{(t-L,t)}^{i}\big)\Big)\in\mathbb{R}^{r},$ (1)
@@
[…]
for all $i\in\{1,\dots,M\}$, and $t\in\{1,\dots,T\}$.
@@
Other architectures are explored in Appendix C.
Unless noted otherwise, we use LongConv Fu et al. (2023) for the sequence layer, two-layer FFNs with dropout for $\phi$, $\rho$, and $\psi$, and chunk size L = 3. MHA-Seq differs only by using five attention heads. Full hyperparameters are in Appendix D.
Proposition 1 highlights the improved time complexity of the forward pass of the Set-Sequence model compared with other approaches to model the cross-section. The proofs of this and other results in this section are given in Appendix A.
Proposition 1 (Forward time for one cross-sectional layer).
Let $M$ be the number of units, $d$ the number of features per unit, and $T$ the sequence length. At each time step $t\in\{1,\dots,T\}$ we hold $M$ vectors $x^{(i)}_{t}\in\mathbb{R}^{d}$. Let $C_{\mathrm{seq}}(T,w)$ denote the cost of one temporal layer on a length-$T$ sequence of width $w$. The forward-time complexity of one cross-sectional layer is:
Method Time Complexity
Set–Seq $\Theta(TMd)+\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$
MHA–Seq $\Theta\!\big(T[M^{2}d+Md^{2}]\big)+\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$
Naïve MHA–Seq $\Theta\!\big(T(Md)^{2}\big)+\Theta\!\big(M\,C_{\mathrm{seq}}(T,d)\big)$
Full-Stacked Temporal $\Theta\!\big(C_{\mathrm{seq}}(T,Md)\big)$
MHA–Seq applies cross-sectional self-attention across the $M$ units, whereas the Naïve MHA–Seq instead treats each feature as a token ($N=Md$). Full-Stacked Temporal applies a single joint temporal model directly to the concatenated $Md$-dimensional features.
Expressivity under exchangeability
When the cross-section at each time is exchangeable—i.e., predictions depend on the multiset of unit features rather than their identities—using a permutation-invariant set summary does not sacrifice expressivity for continuous targets on compact domains. The result below formalizes this: pooled monomial features up to a sufficiently large degree $k$ suffice to approximate any continuous permutation-invariant cross-sectional function.
Proposition 2 (Expressivity of Set Module).
Let $K\subset\mathbb{R}^{d}$ be compact and fix $M\in\mathbb{N}$. Let $G:K^{M}\to\mathbb{R}^{q}$ be continuous and permutation-invariant, i.e., $G(x^{(\pi(1))},\dots,x^{(\pi(M))})=G(x^{(1)},\dots,x^{(M)})$ for every permutation $\pi$. Then for every $\varepsilon>0$ there exist an integer $k=k(\varepsilon,G,K,d,M)\geq 0$, a feature map $\phi_{k}:K\to\mathbb{R}^{\binom{d+k}{k}}$ containing all monomials of total degree $\leq k$, and a continuous map $\rho:\mathbb{R}^{\binom{d+k}{k}}\to\mathbb{R}^{q}$ such that
$\sup_{(x^{(1)},\dots,x^{(M)})\in K^{M}}\Big\|\rho\!\Big(\tfrac{1}{M}\sum_{i=1}^{M}\phi_{k}(x^{(i)})\Big)-G(x^{(1)},\dots,x^{(M)})\Big\|_{\infty}\;<\;\varepsilon.$
Equivalently, the set module $F=\rho\!\big(\tfrac{1}{M}\sum_{i=1}^{M}\phi_{k}(x^{(i)})\big)$ uniformly approximates $G$ on $K^{M}$.
4 Task Selection and Baselines
@@
Common multivariate-forecasting benchmarks—ETTh1, ETTm1, Exchange, and Traffic (see details in Zhou et al. (2021))—are ill-suited for our study: their unit counts are small enough that a full joint model remains tractable or they lack a clear exchangeable-unit structure. We therefore evaluate on three datasets—a synthetic contagion setting, an equity-portfolio regression task, and a mortgage-risk classification task—each defined over exchangeable units. They contain 1000, 500, and 2500 units with 4, 79, and 52 features per unit, giving cross-sectional dimensions of 4000, 39500, and 130000 that meaningfully stress models built for high-dimensional cross-sections.
@@
We benchmark against five widely adopted sequence architectures— Transformer Vaswani et al. (2017), S4 Gu et al. (2022), H3 Fu et al. (2022), Hyena Poli et al. (2023), and LongConv Fu et al. (2023). Together they span attention-, state-space-, and convolution-based methodologies, each of which our Set-Sequence layer can augment. We treat them as task-agnostic baselines on all three datasets. Because these models share a common backbone structure, we hold input/output dimensions, hidden size, learning-rate schedule, and other hyperparameters fixed, so the only variation lies in the sequence layer (and whether a Set module is present). This yields a fair comparison across models. Each task section also reports results for the strongest task-specific domain baseline. Hyperparameters for all models are in Appendix D.
@@ 5 Synthetic Task
We examine the Set-Sequence model’s performance on a synthetic task mimicking loan default prediction, where contagion is a key latent factor Anenberg & Kung (2014); Azizpour et al. (2018); Towe & Lawley (2013). The goal is to predict the next-step transition probabilities for each unit.
We simulate 1000 exchangeable units, each with binary feature $x\in\{0,1\}$ over 100 time steps. Units transition between three states (state 3 is absorbing/default) with a transition matrix proportional to:
@@
The default rate evolves as $\lambda_{x,t+1}=\beta\lambda_{x,t}+\alpha N_{x,t}$, where $N_{x,t}$ is the fraction of type-$x$ units defaulting at time $t$. Parameters $\mu=0.001$, $\alpha=4$, and $\beta=0.5$ yield an average default rate of $\sim$1%. The model captures contagion via $\lambda_{t}=(\lambda_{0,t},\lambda_{1,t})$, where defaults increase future default risk. The core challenge is to model the joint dependencies created by the latent factor $\lambda_{t}$, which is itself a function of each unit’s state.
@@ 5.1 Comparison with Sequence Models
Table 1: Backbone-agnostic evaluation of adding Set–Seq. For each backbone we report KL (↓) and AUC (↑) in three modes: Joint (concatenate all unit features and model a single sequence over the full cross-section), Single (model each unit independently), and Set–Seq (set summary + per unit backbone; values in bold). The rightmost column gives the multiplicative KL reduction of Set–Seq relative to Single.
@@
LongConv 0.037 0.681 0.0018 0.757 0.00018 0.802 10.2
S4 0.038 0.676 0.0016 0.758 0.00019 0.803 8.1
H3 0.040 0.675 0.0017 0.751 0.00039 0.795 4.4
Transformer 0.036 0.506 0.0016 0.758 0.00021 0.801 7.6
Hyena 0.036 0.702 0.0017 0.760 0.00019 0.802 9.1
We compare the Set-Sequence model to the sequence baselines from Section 4 using AUC for the rare default transition and KL divergence between true and predicted transition probabilities. Table 1 reports KL and AUC for each backbone in three modes: Joint (a single sequence over the full cross-section), Single (one unit at a time), and Set–Seq (our set-summary + sequence backbone). Joint models underperform, highlighting the difficulty of modeling large cross-sections with a sparse joint signal. Moving from Joint to Single recovers substantial accuracy, and adding the Set component yields the best results across all backbones: KL improves by $4.4\times$ (H3) to $10.2\times$ (LongConv), with AUC gains of about +0.04+0.04 in each case, indicating that the benefit of the Set summary is robust to the choice of sequence backbone. In later experiments we use the LongConv sequence model backbone.
@@
Seq. Model Set Model Seq. Len.
KL(true|predicted)
AUC
Rel.
Epoch Time
Rel.
Mem
@@
Model Ablations
Table 2 reports the results of ablation study, varying the cross-sectional component (None, Set, or MHA) and sequence length. The largest performance gain comes from adding the Set component. While replacing it with Multi-Head Attention (MHA) further improves accuracy, it incurs a $3.3\times$ training time and $3.5\times$ memory increase due to its quadratic scaling.